2r-75=-5
7r+2s=9
solve with elimination method

Answer by checkley77(12844)

2r-75=-5
2R=-5+75
2R=70
R=70/2
R=35 ANS.
7r+2s=9
7*35+2S=9
245+2S=9
2S=9-245
2S=-236
S=-236/2
S=-118 ANS.
PROOF:
7*35+2*-118=9
245-236=9
9=9